Systems Staff Meeting Notes 8/28/00

Systems Staff Meeting Notes
August 28, 2000

Present: Rich Bennett, Cynthia Bowker, Will Chaney, Debra Harris, Winston Harris, Martha Hruska, Bobbie Meng, Suzy Shaw, Hilton Soto, Geoffrey Stewart, Ying Tang
  1. Winston visited the Dakota Winery over the weekend. Not recommended.
  2. Restaurant review: Grillmasters - great; Ballyhoo - not impressive.
  3. Labor Day Weekend - no Systems meeting next Monday or the following (9/11).
  4. MSNE Training - Starts Saturday, Sept. 2 and goes through the following five Saturdays. Suzy, Will & Bobbie are scheduled to attend. There will be four mandatory tests and two electives. Each test costs $100. IFAS is paying for their staff. Unknown if the libraries have enough money in the staff development or professional development funds to cover. Martha will figure it out. Suzy, Will & Bobbie will be taking one day of comp time per week to make up for the worked Saturday.
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. No more information on money for the ordering of new equipment. Many machines need upgrades and/or replacement. The Systems student machine is one of them. Over 200 machines will be going out of warranty - the 133's. Many are in Resource Services. Some are public machines. Also need to order workstations for new hires. Martha will try to locate the money so we can start ordering by early next week.
  2. VISOR distribution list - Martha is working on it.
  3. HP 4500 Color Printer - is currently in the vault - was originally ordered for Gifts & Exchange, but they don't want it. Will will see about relocating it to DLC and replacing their Mac color printer.
  4. Cynthia's machine arrived last week - it's up and running in room 18A. She is officially working in Systems in the morning.
  5. Need to order eight more switches to populate the closets in West and East North Tower (once renovation is finished).
  6. Dell online support - Web based support did not respond for days on a CD ROM replacement. Then via phone they said they would send a tech support person with the drive to do the actual replacement.
  1. Liaisons will meet next Thursday, Sept. 7 - with Will at the helm.
  1. Wiring for North Tower is still on hold; Digital will be doing the punch down & wiring job. Facilities has them on hold while they wait for baseboards. Rob Roberts says the 2nd floor is done, but Tim from Digital says no. Brad has to drill a hole through a cement wall to pull wires for the 3rd floor. The project is slated to be complete by the middle of September. Doesn't seem likely. Tim is trying to finish up this week if the baseboard supplies show up.
  2. Security Wiring in West - Currently underway - changing to keyless doors, adding cameras, etc. Martha is going to encourage Facilities and other departments' attendance at MMM to keep the middle managers informed of current projects.
  3. Old phone wiring is being removed from West closets.
  4. MS VM patch - there have been a few questions about the instructions Will sent out. Some versions required 2 separate patches (3309 then 3314). It is one of the most complicated patches yet.
  5. VISIO - no new information - Suzy has been in the process of moving out of her old office.
  6. Wireless - still waiting on Brad; money is available. Need to order one more antennae once the project is closer to being finished.
  7. Xerox printer issue - Barbara Oliver called reporting that Marvin Crabb was having trouble with print release stations. Will removed the NLIClear program from the Smathers PS server and the difficulty cleared up. We'll try to put it back on and increase the timeouts. The program seems to prevent the printers from connecting to the server after its been in place for awhile. In the long run, if the conversion to NT goes as planned, we won't have this problem. A new patch for NT should be out in October. The multiple-user sign on is unique to us. An NT version currently only exists with single user sign-on. Xerox took their back-up server - they will install the beta version on it & return it in October. The Music Library will be the test site. If the patch fails, we may have to go to individual sign ons. We really need to formulate a back-up plan in the event that we have to switch to the single sign-ons.
Software Projects Status
  1. Student-Patron updates ran late (Saturday) - this gave David Hickey some problems. There will be one more run before the end of the term. Tigert's IS Department called to verify the problem with 9, 10 & 11 faculty dropping out and those on sabbatical.
  2. DLC permission problems on NT9 - Winston will work on it this evening. He'll wipe all current permissions out & start over with the requested changes.
  3. OPS Tracking changes were made for Cathy Mook.
  4. Personnel, Facilities - projects are currently static.
  5. Hilton - working on Library Stats, film log, LASSIE.
  6. Ying - working on DLC projects.
  7. Cynthia - working on Web version of PAULA.
  8. Martha has inherited Library Statistics due to Carolyn's departure. She is trying to get online reporting on a quarterly basis. Remote entry from Health and Law is going.
DP Coordinator
  1. Students - Andy Zahn's hours will be 10:30 AM-1:00 PM this semester.
  2. When updating or closing trouble tickets, please correspond to SYSHELP rather than Debra's personal email address.
  3. Outlook /laptop reservations - Suzy will get together with Debra to review setup and procedures.
  4. The laptop that should be made available for checkout with the LEC Projector for East needs to be setup and logged in for checkout. It is currently located in the front vault; it needs to be relocated to the tape vault when ready.
  1. Labor Day holiday - September 4.
  2. Bobbie out September 5.
  3. Martha will be out September 11, 26 & 27.
  4. LMS Conference in Orlando - Oct. 2-4 (Martha will be out).

Back to the Systems Homepage

Last updated September 18, 2000
by Debra Harris